OpenStudy (anonymous):

A 2-ft wide circular track for a camera dolly is set up for a movie scene. The two rails of the track form concentric circles. The radius of the inner circle is 67 ft. How much farther does a wheel on the outer rail travel than a wheel on the inner rail of the track in one turn?

OpenStudy (dan815):

Hi :)

OpenStudy (alexandervonhumboldt2):

QH question lol

OpenStudy (anonymous):

this is a qh question?

OpenStudy (dan815):



OpenStudy (dan815):

Okay since the inner circle radius is 67, and the track is 2 feed wide, the outter circle radius is 69, now you need to know the total length of both of these circles

OpenStudy (dan815):

As they are asking you 'how much further' is the travellling distance on the outer track as opposed to the inner track, they are asking you, how much bigger is the 'circumference' of the outer circle compared to the inner circle

OpenStudy (dan815):

\[Circumference of Circle = C=2\pi r\\ C=2 \pi r\\\] let C1 be the circumference of the smaller circle let C2 be the circumference of the bigger circle \[ C1=2* \pi * 67\\ C2=2* \pi * 69\] What is C1 and C2, and what is the difference between these 2 circumferences? Can you figure it out from here?

OpenStudy (anonymous):

oh wow that makes more sense thanks!

OpenStudy (dan815):

okay so.. the difference is the distance travelled on the bigger circumference vs the smaller circumference is \[C2-C1= 2\pi 69- 2 \pi 67=2\pi (69-67) = 4 \pi\]

OpenStudy (anonymous):

ok

OpenStudy (nincompoop):

we are going to need a unit since the question is how far and it has to be consistent with the provided measurement - in feet

OpenStudy (nincompoop):

do you need further help or are you satisfied with the information you were given so far?